Complete Computer (CC) is a retailer of computer equipment in Minneapolis with four retail outlets. Currently...

  1. Complete Computer (CC) is a retailer of computer equipment in Minneapolis with four retail outlets. Currently each outlet manages its ordering independently. Demand at each retail outlet averages 4179 units per week. Each unit costs $200, and CC has a holding cost of 20% per annum. The fixed cost of each order (administrative plus transportation) is $892. Assume 50 weeks in a year.

a)If Item X is ordered whenever the inventory hits zero, and the order quantity is 274 each time an order is placed, what is the annual fixed ordering cost that Complete Computer incurs in one year?

b)If Item X is ordered whenever the inventory hits zero, and the order quantity is 354 each time an order is placed, what is the annual holding cos that Complete Computer incurs in one year? (Note that this is the holding cost for all units of Item X, not just a single unit of Item X).

(please give process)
Answer #1

Demand, d = 4179 units per week

Annual demand, D = 4179*50 = 208950

Holding cost, h = 200*20% = $ 40

Ordering cost, K = 892

a)

Annual fixed orderiing cost at each outlet = (D/q)*K = (208950/274)*892 = 680231.4

Annual fixed ordering cost across the four outlets = 680231.4*4 = $ 2,720,925.6

b)

Annual holding cost at each outlet = (q/2)*h = (354/2)*40 = $ 7080

Annual holding cost across the four outlets = 7080*4 = $ 28,320